an electric force 3.24x10^-4 C exerts an attractive force of 12 N on a second charge that is 0.040 m away. what is the magnitude of the second charge?


1
Expert's answer
2021-05-27T10:07:06-0400

According to Coulomb's law:


"F=k\frac{q_1q_2}{r^2},\\\space\\
q_2=\frac{r^2F}{kq_1}=\frac{0.04^2·12}{(9·10^9)(3.24·10^{-4})}=6.58·10^{-9}\text{ C}."